How to prepare 60 ml of 10% (v/v) H2SO4 solution from Concentrated Sulphuric acid

Grade:10

1 Answers

Arun
6 years ago
Dear Kumar
 
Let the volume of H2SO4 is x
Hence water is 60-x
Now
x*100/(60-x) = 10
10x = 60 -x
11x = 60
x = 60/11 ml = 5.454 ml
 
Thus take 5.454 ml of H2SO4 and 54.546 ml of water
 
Regards
Arun (askIITians forum expert)